# SemioLLM: Evaluating Large Language Models for Diagnostic Reasoning from Unstructured Clinical Narratives in Epilepsy

### v2

@@ 5 Acknowledgments
This work was supported by the Else Kröner Fresenius Foundation, Kolleg Clinbrain: Artificial Intelligence for Clinical Brain Research, the Clinician Scientist program of the Medical Faculty Tübingen funded by Deutsche Forschungsgemeinschaft (DFG, 493665037), the Priority Programme SPP 2241 - PN 520287829 (DFG), the Machine Learning Cluster of Excellence under Germany’s Excellence Strategy – EXC number 2064/1 PN 390727645, the Tübingen AI Center, the Collaborative Research Center 1233 ”Robust Vision” funded by the DFG and ERC (853489-DEXIM). M.D. is a member of the International Max Planck Research School for Intelligent Systems Tübingen (IMPRS-IS). We thank Matthias Tangemann and Elisa Nguyen for comments on the manuscript.
@@ 6 Author Contribution
MD: Conceptualization, Methodology, Validation, Formal Analysis, Data Curation, Reasoning Study Design and Analysis, Visualization, Writing - Original Draft, Writing - Review and Editing. MJP: Methodology (design of clinical study form); Formal Analysis (compilation and evaluation of clinical responses and model source citations), Writing - Review and Editing. ZA: Supervision, Funding Acquisition SL: Conceptualization, Methodology, Clinical evaluation, Project Administration, Supervision, Funding Acquisition, Writing - Review and Editing.

### v3

@@ 4 Methods, Data and Code Availability
The source code to reproduce the findings and figures are present at https://github.com/liebelab/semiollm
@@ 5 Acknowledgments
This work was supported by the Else Kröner Fresenius Foundation, Kolleg Clinbrain: Artificial Intelligence for Clinical Brain Research, the Clinician Scientist program of the Medical Faculty Tübingen funded by Deutsche Forschungsgemeinschaft (DFG, 493665037), the Priority Programme SPP 2241 - PN 520287829 (DFG), the Machine Learning Cluster of Excellence under Germany’s Excellence Strategy – EXC number 2064/1 PN 390727645, the Tübingen AI Center, the Collaborative Research Center 1233 “Robust Vision” funded by the DFG and ERC (853489-DEXIM). M.D. is a member of the International Max Planck Research School for Intelligent Systems Tübingen (IMPRS-IS). The authors thank Prof. Dr. Jakob H. Macke and his lab for discussion and feedback; Almut Sophia Koepke, Shyamgopal Karthik, Matthias Tangemann and Elisa Nguyen for comments on the manuscript.
@@ 6 Author Contribution
MD: Conceptualization, Methodology, Validation, Formal Analysis, Data Curation, Reasoning Study Design and Analysis, Visualization, Writing - Original Draft, Writing - Review and Editing. MJP: Methodology (design of clinical study form); Formal Analysis (compilation and evaluation of clinical responses and model source citations), Writing - Review and Editing. FP: Clinical evaluation, Writing - Review. ZA: Supervision, Funding Acquisition. SL: Conceptualization, Methodology, Clinical evaluation, Project Administration, Supervision, Funding Acquisition, Writing - Review and Editing.